OpenStudy (anonymous):

Please help I will fan and give medal:

OpenStudy (anonymous):

Divide: \[\frac{ x^{2}+11x+30 }{ x^2{-25} } \div \frac{ x+6 }{ x-6 }\]

rishavraj (rishavraj):

\[x^2 + 11x + 30 = (x + 5)(x + 6)\] \[x^2 - 25 = x^2 - 5^2 = (x + 5) (x - 5)\]

rishavraj (rishavraj):

so now it is \[\frac{ (x + 5)(x + 6) }{ (x + 5)(x - 5) } \div \frac{ (x + 6) }{ (x - 6) }\]

OpenStudy (anonymous):

oh wow

OpenStudy (anonymous):

ok so now i cancel out and get \[\frac{ (x+6) }{ (x-5) }\div \frac{ (x+6) }{ (x-6) }\]

OpenStudy (anonymous):

@rishavraj ^^^^

rishavraj (rishavraj):

yup.. u know \[\frac{ a }{ b } \div \frac{ c }{ d } = \frac{ ad }{ bc } \]

OpenStudy (anonymous):

oh cross multiply

rishavraj (rishavraj):

hmmmmm

OpenStudy (anonymous):

so my final answer after canceling out would be: \[\frac{ (x-5) }{ (x-6) } ?\]

rishavraj (rishavraj):

\[\frac{ (x - 6) }{ (x - 5) }\]

OpenStudy (anonymous):

how?

rishavraj (rishavraj):

\[\frac{ (x + 6) }{ (x - 5) } \div \frac{ (x + 6) }{ (x - 6) }\] = \[\frac{ (x + 6)(x - 6) }{ (x - 5)(x + 6) }\]

OpenStudy (anonymous):

ok so my final answer would be (x-6)/(x-5) ?

OpenStudy (anonymous):

@rishavraj final answer: \[\frac{ (x-6) }{ (x-5) }\]

rishavraj (rishavraj):

if u r supposed to simplify it then YES

OpenStudy (anonymous):

Thanks so much

rishavraj (rishavraj):

yw:PP